OpenStudy (anonymous):

Anne and Nancy use a metal alloy that is 20% copper to make jewelry. How many ounces of a 14% alloy must be mixed with a 23% alloy to form 69 ounces of the desired alloy?

OpenStudy (kropot72):

The amount of copper required in the final 69 ounces of alloy = 69 * (20/100) = 13.8 ounces. Let the required quantity of 14% alloy = a Let the required quantity of 23% alloy = b a + b = 69 a = 69 - b 0.14a + 0.23b = 13.8 0.14(69 - b) + 0.23b = 13.8 9.66 - 0.14b + 0.23b = 13.8 0.09b = 4.4 Can you find b now?